The sum of the digits of a two-digit number is 15. If the digits are reversed, the new number is only 9 more than the original number. What was the original number?
Answer by jorel1380(3719)

m+n=15
m=15-n
10m+n=10n+m+9
10(15-n)+n=10n+15-n+9
150-9n=9n+24
18n=126
n=7
m=8
The original number was 78..